OpenStudy (anonymous):

Translate this problem to an equation and solve. The product of 2 positive consecutive even integers is 48. What is the smaller number?

hartnn (hartnn):

you know the meaning of "consecutive" ?

OpenStudy (anonymous):

Product means multiply. Let the first even number be x and the consecutive even number be x + 2.

hartnn (hartnn):

you know how to solve quadratic equations ?

OpenStudy (anonymous):

no i dont know how to solve it

hartnn (hartnn):

did u get the equation ?

OpenStudy (anonymous):

no i dont get how to do it

OpenStudy (radar):

Do you get how AnimEfReaK assigned the variables, (x) and (x+2)?

OpenStudy (radar):

Further did you get how product indicates multiplication?

OpenStudy (radar):

Which means x(x+2)=48 or x^2 + 2x = 48.

OpenStudy (anonymous):

first integer = x 2nd integer = x + 2 x(x + 2) = 48 -- distribute x^2 + 2x = 48 x^2 + 2x - 48 = 0 quadratic formula : (-b +- sq r b^2 - 4ac) 2a a = 1, b = 2, and c = -48 now we sub (-2 +- sq r 2^2 - 4(1)(-48)) / 2(1) (-2 +- sq r 4 + 192) / 2 (-2 +- sq r 196)/2 (-2 +- 14)/2 (-2 + 14) /2 = 12/2 = 6 (-2 - 14)/2 = -16/2 = -8 smaller number is -8

OpenStudy (radar):

\[x ^{2}+2x -48=0\]This factors nicely which is a method of solving a quadratic. (x-6)(x+8) are the factors. Only one results in a positive integer x-6 = 0 x=6 x+2 = 8